OpenStudy (anonymous):

-3 1/5 + (-6 1/2) A.) -3 2/7 B.) -3 7/10 C.) -9 2/7 D.) -9 7/10

OpenStudy (tkhunny):

= (-3 - 6) - (1/5 + 1/2) = -9 - (2/10 + 5/10) = -9 - 7/10 = -9 7/10 It's a little annoying to keep track of that.

OpenStudy (anonymous):

thank you

OpenStudy (tkhunny):

Not everyone is willing to work with mixed fractions in this way. It's okay, really!